An electrically neutral model airplane is flying in a horizontal cirele on a 3.5-m guideline, which is nearly parallel to the ground. The line breaks when the kinetic energy of the plane is so J Recansider the same stuation, excent that now there is a point charge of tg on the plane and a point charge of g at the other end of the guideline. In this case, the line breaks when the kinetic energy of the plane is 52.D 3. Find the magnitude of the charges. 2.011e-4
Answer #1

We know that,

KE = (1/2)mv^2

F = mv^2 /r

Solving both, F = 2KE/r

F = 2*50/3.5 = 28.57 N

The line breaks when the tension is 28.57 N

F' = 2*52/3.5 = 29.71 N

Now, Resultant force F'

F' = F + Fa

Where, Fa is attraction force. So it is negative

Fa = 29.71 - 28.57 = 1.14 N

By using Columbs law,

F = Kq1*q2/d^2

- 1.14 = 9*10^9*(+q)*(-q)/(3.5)^2

q^2 = 1.14*(3.5)^2 /(9*10^9)

q = 4*10^-5 C